# A complete classification of the zero-dimensional homogeneous spaces under determinacy

### v2

@@ 2.2. Determinacy and nice topological pointclasses
For more extensive introductions to the topic of games, we refer the reader to [Ke, Section 20] or [Je, Chapter 33]. Here, we only want to give the definitions of strategy and determinacy. Given a set $\Omega$ and $S\subseteq\Omega^{\omega}$, a play of the game $\mathsf{G}=\mathsf{G}(\Omega,S)$ is described by the diagram
I $x_{0}$ $x_{1}$cdots
II $y_{0}$ $y_{1}$cdots,
@@
A strategy for Player I on $\Omega$ is a monotone $\sigma:\Omega^{<\omega}\longrightarrow\Omega^{<\omega}$ such that $|\sigma(s)|=|s|+1$ for every $s\in\Omega^{<\omega}$. We will say that a strategy $\sigma$ for Player I on $\Omega$ is winning in $\mathsf{G}$ if setting $x=f_{\sigma}(y)$ yields a play in which Player I wins for every $y\in\Omega^{\omega}$, where $x=(x_{0},x_{1},\ldots)$ and $y=(y_{0},y_{1},\ldots)$. Similarly, a strategy for Player II on $\Omega$ is a monotone $\tau:\Omega^{<\omega}\longrightarrow\Omega^{<\omega}$ such that $|\tau(s)|=|s|$ for all $s\in\Omega^{<\omega}$, and $\tau$ is winning in $\mathsf{G}$ if setting $y=f_{\tau}(x)$ yields a play in which Player II wins for every $x\in\Omega^{\omega}$. We will say that the game $\mathsf{G}$ (or simply the set $S$) is determined if (exactly) one of the players has a winning strategy. For the purposes of this article, only the case in which $\Omega$ is countable will be relevant. Given $\mathbf{\Sigma}\subseteq\mathcal{P}(\omega^{\omega})$, we will write $\mathsf{Det}(\mathbf{\Sigma})$ to mean that every element of $\mathbf{\Sigma}$ is determined. The assumption $\mathsf{Det}(\mathcal{P}(\omega^{\omega}))$ is known as the Axiom of Determinacy ($\mathsf{AD}$). The assumption $\mathsf{Det}(\bigcup_{1\leq n<\omega}\mathbf{\Sigma}^{1}_{n}(\omega^{\omega}))$ is known as the axiom of Projective Determinacy ($\mathsf{PD}$).
@@
As the determinacy assumptions get stronger, the set-theoretic universe becomes more regular. It follows that $\mathsf{AC}$ is incompatible with $\mathsf{AD}$, since $\mathsf{AC}$ enables one to construct pathological sets (see for example [Je, Lemma 33.1]). For this reason, as we mentioned at the beginning of the article, we will be working in $\mathsf{ZF}+\mathsf{DC}$333 As eloquently put by Kanamori, while comparing $\mathsf{DC}$ to a weaker axiom, “it soon became de rigueur to assume the stronger $\mathsf{DC}$ in much of the investigation of $\mathsf{AD}$” (see [Ka, page 378]). It is a famous theorem of Martin that $\mathsf{Det}(\mathsf{Borel}(\omega^{\omega}))$ holds in $\mathsf{ZF}+\mathsf{AC}_{\omega}$ (this was originally proved in [Ma1], but see also [Ma2, Remark (2) on page 307]). However, Harrington showed that $\mathsf{Det}(\mathbf{\Sigma}^{1}_{1}(\omega^{\omega}))$ has large cardinal strength (see [Har]). For the consistency of $\mathsf{ZF}+\mathsf{DC}+\mathsf{AD}$, see [Ne] and [Ka, Proposition 11.13].
Most results that use determinacy assumptions follow a consistent pattern: they will hold for Borel sets without any additional assumptions (by Martin’s theorem), for projective sets under $\mathsf{PD}$, and for all sets under $\mathsf{AD}$. The following is [CMM2, Definition 3.1], and it provides a convenient expositional tool to simultaneously state the various versions of our results.444 Notice that the term “function” in the following definition is an abuse of terminology, as each topological pointclass is a proper class. Therefore, every theorem in this paper that mentions these pointclasses is strictly speaking an infinite scheme (one theorem for each topological pointclass). This is what we mean when we speak about “level-by-level” results. Given a set $Z$ and $\mathbf{\Sigma}\subseteq\mathcal{P}(Z)$, we will denote by $\mathsf{b}\mathbf{\Sigma}$ the smallest boolean subalgebra of $\mathcal{P}(Z)$ that contains $\mathbf{\Sigma}$.
@@
For the purposes of this paper, the following are the intended examples of nice topological pointclasses.555 For a more detailed discussion, see [CMM2, Section 3]. Recall that $\mathbf{\Sigma}^{1}_{n}(Z)$ for $1\leq n<\omega$ can be defined for an arbitrary space $Z$ by declaring $A\in\mathbf{\Sigma}^{1}_{n}(Z)$ if there exists a Polish space $W$ and an embedding $j:Z\longrightarrow W$ such that $j[A]=\widetilde{A}\cap j[Z]$ for some $\widetilde{A}\in\mathbf{\Sigma}^{1}_{n}(Z)$.
@@
We conclude this section by giving precise formulations of well-known results concerning the Baire property under suitable determinacy assumption. We remark that Corollary 2.8 will be needed in the proof of Theorem 9.9. Recall that a subset $A$ of a space $Z$ has the Baire property if there exists an open subset $U$ of $Z$ such that $A\Delta U$ is a meager subset of $Z$.
@@
and observe that $\varphi:\Omega^{\omega}\longrightarrow\omega^{\omega}$ is continuous. In order to prove that $A$ has the Baire property in $\omega^{\omega}$, by [Ke, Exercises 8.35 and 8.36], it will be enough to show that $\mathsf{G}(\Omega,\varphi^{-1}[(\omega^{\omega}\setminus A)\cup U])$ is determined for every open subset $U$ of $\omega^{\omega}$, where $\mathsf{G}$ is the game introduced at the beginning of this section. This is straightforward to check, using the fact that $\mathbf{\Sigma}$ is a nice topological pointclass and our determinacy assumptions. ∎
@@ 2.3. Filters and semifilters
Throughout this section, we will assume that $\Omega$ is a given countably infinite set. We will say that $\mathcal{X}\subseteq\mathcal{P}(\Omega)$ is closed under finite modifications if $|x\Delta y|<\omega$ and $x\in\mathcal{X}$ implies $y\in\mathcal{X}$ for all $x,y\subseteq\Omega$. A collection $\mathcal{X}$ is closed under finite intersections (respectively closed under finite unions) if $x\cap y\in\mathcal{X}$ (respectively $x\cup y\in\mathcal{X}$) for all $x,y\in\mathcal{X}$. A collection $\mathcal{X}\subseteq\mathcal{P}(\Omega)$ is upward-closed (respectively downward-closed) if $y\supseteq x\in\mathcal{X}$ (respectively $y\subseteq x\in\mathcal{X}$) implies $y\in\mathcal{X}$ for all $x,y\in\mathcal{P}(\Omega)$.
@@
$\left.\begin{array}[]{lcl}&&h_{F}(x)(n)=\left\{\begin{array}[]{ll}1-x(n)&\textrm{if }n\in F,\\ x(n)&\textrm{if }n\in\omega\setminus F.\end{array}\right.\end{array}\right.$
@@
Another notable topological property of filters is given by the following result from [MZ2], which will be useful in Section 10.2.
@@
We conclude this section with some well-known facts about the Baire category of the above combinatorial objects under suitable determinacy assumptions.
@@ 2.4. Homogeneity
A zero-dimensional space $X$ is strongly homogeneous (or h-homogeneous) if every non-empty clopen subspace of $X$ is homeomorphic to $X$.
Familiar examples of strongly homogeneous spaces are given by the rationals $\mathbb{Q}$, the Cantor set $2^{\omega}$, and Baire space $\omega^{\omega}$. This can be seen by using their characterizations (see [Ke, Exercise 7.12, Theorem 7.4 and Theorem 7.7] respectively).
As an unpleasant example, consider the subspace
$X=(\mathbb{R}\times\{0\})\cup\{(q_{n},2^{-n}):n\in\omega\}$
of $\mathbb{R}\times\mathbb{R}$, where =$\mathbb{Q}=\{q_{n}:n\in\omega\}$ is an enumeration. Using [Me1, Theorem 18], one can show that every non-empty clopen subspace of $X^{\omega}$ is homeomorphic to $X^{\omega}$. On the other hand, the space $X^{\omega}$ is not homogeneous because it is zero-dimensional at some points but not at others. So, if we dropped the zero-dimensionality requirement from Definition 2.13, there would be strongly homogeneous spaces that are not homogeneous, which is highly undesirable.
In fact, even when confined to the zero-dimensional realm, the relationship between homogeneity and strong homogeneity remains rather delicate. It is a folklore result that every strongly homogeneous space is homogeneous (see Corollary 2.16). Furthermore, the discrete spaces of size at least two and $\omega\times 2^{\omega}$ are obviously homogeneous but not strongly homogeneous. However, as Proposition 2.14 and Theorem 9.19 will show, it is consistent that these are the only zero-dimensional counterexamples (this was first proved in [CMM1]).
@@
The following ingenious result is an important tool for constructing homogeneous spaces, and it first appeared as [vM2, Lemma 2.1].666 This result was later generalized by Medvedev (see [Mv, Theorem 3]). In fact, the proof given here is a simplified version of Medvedev’s proof. Corollary 2.17 is essentially the same as [MvMZ, Lemma 2]. Given a space $X$ and $x\in X$, recall that a local base for $X$ at $x$ is a collection $\mathcal{B}$ consisting of open neighborhoods of $x$ such that for every open neighborhood $U$ of $x$ there exists $V\in\mathcal{B}$ such that $V\subseteq U$.
@@
Proof.
Assume without loss of generality that $U_{0}\cap V_{0}=\varnothing$. Fix homeomorphisms $f_{n}:U_{n}\longrightarrow V_{n}$ for $n\in\omega$. If $f_{n}(x)=y$ for some $n$, then it would be clear how to construct the desired homeomorphism. So assume that $f_{n}(x)\neq y$ for each $n$. This makes it possible to also assume that $f_{n}[U_{n+1}]\cap V_{n+1}=\varnothing$ for each $n$. We will construct homeomorphisms $h_{n}:U_{n}\setminus U_{n+1}\longrightarrow V_{n}\setminus V_{n+1}$ for $n\in\omega$. In the end, set
$\left.\begin{array}[]{lcl}&&h(z)=\left\{\begin{array}[]{ll}z&\textrm{if }z\in X\setminus(U_{0}\cup V_{0}),\\ h_{n}(z)&\textrm{if }U_{n}\setminus U_{n+1},\\ h_{n}^{-1}(z)&\textrm{if }V_{n}\setminus V_{n+1},\\ y&\textrm{if }z=x,\\ x&\textrm{if }z=y.\end{array}\right.\end{array}\right.$
for $z\in X$. It is straightforward to check that $h$ will be as desired.
So pick $n\in\omega$. Recall that $f_{n}[U_{n+1}]\cap V_{n+1}=\varnothing$, hence $U_{n+1}\cap f_{n}^{-1}[V_{n+1}]=\varnothing$. Observe that $f_{n}[U_{n+1}\cup f_{n}^{-1}[V_{n+1}]]=f_{n}[U_{n+1}]\cup V_{n+1}$, therefore
$f_{n}[U_{n}\setminus(U_{n+1}\cup f_{n}^{-1}[V_{n+1}])]=V_{n}\setminus(f_{n}[U_{n+1}]\cup V_{n+1}).$
Furthermore, since
$f_{n}^{-1}[V_{n+1}]\approx V_{n+1}\approx U_{n+1}\approx f_{n}[U_{n+1}],$
it is possible to fix a homeomorphism $g_{n}:f_{n}^{-1}[V_{n+1}]\longrightarrow f_{n}[U_{n+1}]$. Finally, given $z\in U_{n}\setminus U_{n+1}$, define
$\left.\begin{array}[]{lcl}&&h_{n}(z)=\left\{\begin{array}[]{ll}f_{n}(z)&\textrm{if }z\in U_{n}\setminus(U_{n+1}\cup f_{n}^{-1}[V_{n+1}]),\\ g_{n}(z)&\textrm{if }z\in f_{n}^{-1}[V_{n+1}].\end{array}\right.\end{array}\right.$
It is clear that $h_{n}:U_{n}\setminus U_{n+1}\longrightarrow V_{n}\setminus V_{n+1}$ is a homeomorphism, as desired. ∎
@@
Every zero-dimensional strongly homogeneous space is homogeneous.
Corollary 2.17.
@@
So pick $n\in\omega$, and set $F=\{i\in n:x(i)\neq y(i)\}$. Consider the homeomorphism $h_{F}$ defined in Section 2.3, and observe that $h_{F}[\mathsf{N}_{x\upharpoonright n}]=\mathsf{N}_{y\upharpoonright n}$. Using the fact that $\mathcal{X}$ is closed under finite modifications, one sees that $h_{F}[\mathcal{X}]=\mathcal{X}$. It follows that $h_{F}[U_{n}]=V_{n}$, hence $U_{n}\approx V_{n}$. ∎
@@
The following diagram illustrates the connections between the homogeneity-type properties and the combinatorial properties that we have discussed so far. With the possible exception of the implication mentioned by Question 2.20, the implications depicted below (and their obvious consequences) are the only ones that are provable in $\mathsf{ZF}+\mathsf{DC}$ alone. The desired counterexamples are given by Proposition 10.10, by the proof of [CMM1, Theorem 1.2] (which was essentially obtained by van Douwen in [vD]), and by [Me3, Proposition 8.3].777 A few clarifications are in order, so that trivial counterexamples will be avoided. First, we are restricting the attention to zero-dimensional spaces that are not locally compact. Second, by filter (respectively semifilter), we mean homeomorphic to a filter (respectively homeomorphic to a semifilter). Third, regarding the relationship between filters and topological groups, we refer the reader to Section 10.4 for a more nuanced discussion.
${\mathrm{Filter}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$Topological ${\mathrm{Topological\text{ }group}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Semifilter}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$Strongly ${\mathrm{Strongly\text{ }homogeneous}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Homogeneous}}$
@@
We conclude this section with a technical result that will be needed in the proof of Corollary 9.10. It is a slight variation on [Te, Theorem 2.4]. Recall that a $\pi$-base $\mathcal{B}$ for a space is a collection $\mathcal{B}$ consisting of non-empty open sets such that for every non-empty open set $U$ there exists $V\in\mathcal{B}$ such that $V\subseteq U$.
@@
It is easy to realize that if $X$ has an isolated point then $|X|=1$, so assume that $X$ is crowded. If $X$ is compact then $X\approx 2^{\omega}$ by [Ke, Theorem 7.4], so assume that $X$ is non-compact. Fix an infinite partition of $X$ consisting of clopen sets, and injectively enumerate it as $\{X_{n}:n\in\omega\}$.
Pick a non-empty clopen subspace $U$ of $X$, and notice that $U$ contains a clopen subspace that is homeomorphic to $X$. In particular, we can fix an infinite partition of $U$ consisting of clopen sets, and injectively enumerate it as $\{U_{n}:n\in\omega\}$. Start by setting $Y_{0}=\varnothing$, then use our assumptions on $\mathcal{B}$ to recursively construct clopen sets $V_{n}\subseteq U_{n}$ and $Y_{n+1}\subseteq X_{n+1}$ for $n\in\omega$ such that $V_{n}\approx X_{n}\setminus Y_{n}$ and $Y_{n+1}\approx U_{n}\setminus V_{n}$ for each $n$. Finally, fix homeomorphisms $f_{n}:X_{n}\setminus Y_{n}\longrightarrow V_{n}$ and $g_{n}:Y_{n+1}\longrightarrow U_{n}\setminus V_{n}$ for $n\in\omega$, and set $h=\bigcup_{n\in\omega}(f_{n}\cup g_{n})$. It is straightforward to check that $h:X\longrightarrow U$ is a homeomorphism, as desired. ∎
Chapter 3 Wadge theory: fundamental notions and results
The purpose of this chapter is to establish the Wadge-theoretic preliminaries for the remainder of this monograph, mostly referring to [CMM2] for the proofs. The two most fundamental results are Wadge’s Lemma 3.3 and the fact that the Wadge hierarchy is well-founded (see Theorem 3.4). The method of relativization allows one to apply Wadge theory not just in $\omega^{\omega}$ (which is usually the only ambient space considered in the literature), but more generally in zero-dimensional Polish (or even Borel) spaces. We will also discuss the notions of level and expansion, which are best understood together (thanks to the Expansion Theorem 3.25). Finally, we will discuss separated differences, which can be used to give a concrete description of the non-selfdual Wadge classes of level 00. The operation of separated differences will be a crucial technical tool for both the uniqueness and existence results, because of the role that it plays in obtaining the closure properties of Sections 7.2 and 7.5.

### v3

@@ 2.2. Determinacy and nice topological pointclasses
For more extensive introductions to the topic of games, we refer to [Ke, §20] or [Je, Chapter 33]. Here, we only want to give the definitions of strategy and determinacy. Given a set $\Omega$ and $S\subseteq\Omega^{\omega}$, a play of the game $\mathsf{G}=\mathsf{G}(\Omega,S)$ is described by the diagram
I $x_{0}$ $x_{1}$ $\cdots$
II $y_{0}$ $y_{1}$ $\cdots$,
@@
A strategy for Player I on $\Omega$ is a monotone $\sigma:\Omega^{<\omega}\longrightarrow\Omega^{<\omega}$ such that $|\sigma(s)|=|s|+1$ for every $s\in\Omega^{<\omega}$. We will say that a strategy $\sigma$ for Player I on $\Omega$ is winning in $\mathsf{G}$ if setting $x\upharpoonright n+1=\sigma(y\upharpoonright n)$ for $n\in\omega$ yields a play in which Player I wins for every $y\in\Omega^{\omega}$, where $x=(x_{0},x_{1},\ldots)$ and $y=(y_{0},y_{1},\ldots)$. Similarly, a strategy for Player II on $\Omega$ is a monotone $\tau:\Omega^{<\omega}\longrightarrow\Omega^{<\omega}$ such that $|\tau(s)|=|s|$ for all $s\in\Omega^{<\omega}$, and $\tau$ is winning in $\mathsf{G}$ if setting $y\upharpoonright n=\tau(x\upharpoonright n)$ for $n\in\omega$ yields a play in which Player II wins for every $x\in\Omega^{\omega}$. We will say that the game $\mathsf{G}$ (or simply the set $S$) is determined if (exactly) one of the players has a winning strategy. For the purposes of this article, only the case in which $\Omega$ is countable will be relevant. Given $\mathbf{\Sigma}\subseteq\mathcal{P}(\omega^{\omega})$, we will write $\mathsf{Det}(\mathbf{\Sigma})$ to mean that every element of $\mathbf{\Sigma}$ is determined. The assumption $\mathsf{Det}(\mathcal{P}(\omega^{\omega}))$ is known as the Axiom of Determinacy ($\mathsf{AD}$). The assumption $\mathsf{Det}(\bigcup_{1\leq n<\omega}\mathbf{\Sigma}^{1}_{n}(\omega^{\omega}))$ is known as the axiom of Projective Determinacy ($\mathsf{PD}$).
@@
As the determinacy assumptions get stronger, the set-theoretic universe becomes more regular. It follows that $\mathsf{AC}$ is incompatible with $\mathsf{AD}$, since $\mathsf{AC}$ enables one to construct pathological sets (see for example [Je, Lemma 33.1]). For this reason, as we mentioned at the beginning of the article, we will be working in $\mathsf{ZF}+\mathsf{DC}$.444 As eloquently put by Kanamori, while comparing $\mathsf{DC}$ to a weaker axiom, “it soon became de rigueur to assume the stronger $\mathsf{DC}$ in much of the investigation of $\mathsf{AD}$” (see [Ka, page 378]). It is a famous theorem of Martin that $\mathsf{Det}(\mathsf{Borel}(\omega^{\omega}))$ holds in $\mathsf{ZF}+\mathsf{AC}_{\omega}$ (this was originally proved in [Ma1], but see also [Ma2, Remark (2) on page 307]). However, Harrington showed that $\mathsf{Det}(\mathbf{\Sigma}^{1}_{1}(\omega^{\omega}))$ has large cardinal strength (see [Har]). For the consistency of $\mathsf{ZF}+\mathsf{DC}+\mathsf{AD}$, see [Ne] and [Ka, Proposition 11.13].
Most results that use determinacy assumptions follow a consistent pattern: they will hold for Borel sets without any additional assumptions (by Martin’s theorem), for projective sets under $\mathsf{PD}$, and for all sets under $\mathsf{AD}$. The following is [CMM2, Definition 3.1], and it provides a convenient expositional tool to simultaneously state the various versions of our results.555 Notice that the term “function” in the following definition is an abuse of terminology, as each topological pointclass is a proper class. Therefore, every theorem in this paper that mentions these pointclasses is strictly speaking an infinite scheme (one theorem for each topological pointclass). This is what we mean when we speak about “level-by-level” results. Given a set $Z$ and $\mathbf{\Sigma}\subseteq\mathcal{P}(Z)$, we will denote by $\mathsf{b}\mathbf{\Sigma}$ the smallest boolean subalgebra of $\mathcal{P}(Z)$ that contains $\mathbf{\Sigma}$.
@@
For the purposes of this paper, the following are the intended examples of nice topological pointclasses.666 For a more detailed discussion, see [CMM2, §3]. Recall that $\mathbf{\Sigma}^{1}_{n}(Z)$ for $1\leq n<\omega$ can be defined for an arbitrary space $Z$ by declaring $A\in\mathbf{\Sigma}^{1}_{n}(Z)$ if there exists a Polish space $W$ and an embedding $j:Z\longrightarrow W$ such that $j[A]=\widetilde{A}\cap j[Z]$ for some $\widetilde{A}\in\mathbf{\Sigma}^{1}_{n}(Z)$.
@@
We conclude this subsection by giving precise formulations of well-known results concerning the Baire property under suitable determinacy assumption. We remark that Corollary 2.7 will be needed in the proof of Theorem 6.8. Recall that a subset $A$ of a space $Z$ has the Baire property if there exists an open subset $U$ of $Z$ such that $A\Delta U$ is a meager subset of $Z$.
@@
and observe that $\varphi:\Omega^{\omega}\longrightarrow\omega^{\omega}$ is continuous. In order to prove that $A$ has the Baire property in $\omega^{\omega}$, by [Ke, Exercises 8.35 and 8.36], it will be enough to show that $\mathsf{G}(\Omega,\varphi^{-1}[(\omega^{\omega}\setminus A)\cup U])$ is determined for every open subset $U$ of $\omega^{\omega}$, where $\mathsf{G}$ is the game introduced at the beginning of this subsection. This is straightforward to check, using the fact that $\mathbf{\Sigma}$ is a nice topological pointclass and our determinacy assumptions. ∎
@@ 2.3. Filters and semifilters
Throughout this subsection, we will assume that $\Omega$ is a given countably infinite set. We will say that $\mathcal{X}\subseteq\mathcal{P}(\Omega)$ is closed under finite modifications if $|x\Delta y|<\omega$ and $x\in\mathcal{X}$ implies $y\in\mathcal{X}$ for all $x,y\subseteq\Omega$. A collection $\mathcal{X}$ is closed under finite intersections (respectively closed under finite unions) if $x\cap y\in\mathcal{X}$ (respectively $x\cup y\in\mathcal{X}$) for all $x,y\in\mathcal{X}$. A collection $\mathcal{X}\subseteq\mathcal{P}(\Omega)$ is upward-closed (respectively downward-closed) if $y\supseteq x\in\mathcal{X}$ (respectively $y\subseteq x\in\mathcal{X}$) implies $y\in\mathcal{X}$ for all $x,y\in\mathcal{P}(\Omega)$.
@@
[…]
@@
Another notable topological property of filters is given by the following result from [MZ2], which will be useful in §7.2.
@@
We conclude this subsection with some well-known facts about the Baire category of the above combinatorial objects under suitable determinacy assumptions.
@@ 2.4. Homogeneity
A space $X$ is strongly homogeneous (or h-homogeneous) if $X$ is zero-dimensional and every non-empty clopen subspace of $X$ is homeomorphic to $X$.
Familiar examples of strongly homogeneous spaces are given by the rationals $\mathbb{Q}$, the Cantor set $2^{\omega}$, and Baire space $\omega^{\omega}$. This can be seen by using their characterizations (see [Ke, Exercise 7.12, Theorem 7.4 and Theorem 7.7] respectively). The reason for the modifier “strongly” is given by Corollary 2.15, which also explains why we included zero-dimensionality in the definition.777 To see that the assumption of zero-dimensionality is indispensable, consider the subspace $X=(\mathbb{R}\times\{0\})\cup\{(q_{n},2^{-n}):n\in\omega\}$ of $\mathbb{R}\times\mathbb{R}$, where $\mathbb{Q}=\{q_{n}:n\in\omega\}$ is an enumeration. Using [Me1, Theorem 18], one can show that every non-empty clopen subspace of $X^{\omega}$ is homeomorphic to $X^{\omega}$. On the other hand, it is easy to realize that $X^{\omega}$ is not homogeneous.
In fact, even when confined to the zero-dimensional realm, the relationship between homogeneity and strong homogeneity remains rather delicate. The discrete spaces of size at least two and $\omega\times 2^{\omega}$ are obviously homogeneous but not strongly homogeneous. However, as Proposition 2.13 and Theorem 6.18 will show, it is consistent that these are the only zero-dimensional counterexamples (this was first proved in [CMM1]).
@@
The following ingenious result is an important tool for constructing homogeneous spaces, and it first appeared as [vM, Lemma 2.1]. Corollary 2.16 is essentially the same as [MvMZ, Lemma 2].
@@
Every strongly homogeneous space is homogeneous.
Corollary 2.16 (Medini, van Mill, Zdomskyy).
@@
So pick $n\in\omega$, and set $F=\{i\in n:x(i)\neq y(i)\}$. Consider the homeomorphism $h_{F}$ defined in §2.3, and observe that $h_{F}[\mathsf{N}_{x\upharpoonright n}]=\mathsf{N}_{y\upharpoonright n}$. Using the fact that $\mathcal{X}$ is closed under finite modifications, one sees that $h_{F}[\mathcal{X}]=\mathcal{X}$. It follows that $h_{F}[U_{n}]=V_{n}$, hence $U_{n}\approx V_{n}$. ∎
@@
The following diagram illustrates the connections between the homogeneity-type properties and the combinatorial properties that we have discussed so far. With the possible exception of the implication mentioned by Question 2.19, the implications depicted below (and their obvious consequences) are the only ones that are provable in $\mathsf{ZF}+\mathsf{DC}$ alone. The desired counterexamples are given by Proposition 7.10, by the proof of [CMM1, Theorem 1.2] (which was essentially obtained by van Douwen in [vD]), and by [Me3, Proposition 8.3].888 A few clarifications are in order, so that trivial counterexamples will be avoided. First, we are restricting the attention to zero-dimensional spaces that are not locally compact. Second, by filter (respectively semifilter), we mean homeomorphic to a filter (respectively homeomorphic to a semifilter). Third, regarding the relationship between filters and topological groups, we refer to §7.4 for a more nuanced discussion.
${\mathrm{Filter}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Topological\text{ }group}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Semifilter}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Strongly\text{ }homogeneous}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Homogeneous}}$
@@
We conclude this subsection with a technical result that will be needed in the proof of Corollary 6.9. It is a slight variation on [Te, Theorem 2.4].
@@
It is easy to realize that if $X$ has an isolated point then $|X|=1$, so assume that $X$ is crowded. If $X$ is compact then $X\approx 2^{\omega}$ by [Ke, Theorem 7.4]. Otherwise, proceed as in the proof of [Te, Theorem 2.4] (see also [Me1, Appendix A] or [Me2, Appendix B]). ∎
3. Wadge theory: fundamental notions and results